A room is 4 m long and 3 m 50 cm wide. Howe many square metres of carpet is needed to cover the floor of the room?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
l = 4m
b = 3m 50 cm = 3.50 m
Area = l × b = 4 × 3.50
= 14m^2

Thanks